undeadknight26 (undeadknight26):

A little help here? Inverse? F(x) = 7x + 2

undeadknight26 (undeadknight26):

So far i have: F(x) = 7x + 2 Y = 7x + 2 X = 7y + 2 Sqrtx = sqrt7y + 2

OpenStudy (anonymous):

Where are you not getting??

undeadknight26 (undeadknight26):

I do not know where to go from here.

OpenStudy (anonymous):

Why are you taking square root in last step everywhere??

OpenStudy (idealist10):

y=7x+2 x=7y+2 7y=x-2 y=(x-2)/7 Answer: y=(x-2)/7

OpenStudy (anonymous):

See, when you flip x and y, after that very equation, you need to find y in terms of x..

OpenStudy (idealist10):

After you get x=7y+2, you just solve for y in terms of x.

OpenStudy (anonymous):

Your first three steps are okay, don't do last step if you have linear function..

undeadknight26 (undeadknight26):

Thats is the way they taught me @waterineyes. And that makes alot of sense @Idealist10 ! Thank you!

OpenStudy (anonymous):

They?? Who??

OpenStudy (idealist10):

You're welcome!

undeadknight26 (undeadknight26):

Can you help me with a another one? Its related to this!

OpenStudy (anonymous):

No much how many of them you are trying to solve, if you are solving questions of only one type, then I suggest you to do only 2 problems of that type..